OpenStudy (anonymous):

Solve for the roots in the equation below. x^4 + 3x^2 - 4 = 0

hartnn (hartnn):

take \(x^2 =t \) then \(x^4 = (x^2)^2 = t^2\) so you will get a quadratic in 't' can you solve quadratics ?

OpenStudy (anonymous):

no. is it 9 and 16

OpenStudy (cwrw238):

t^2 + 3t - 4 = 0 can you factor this ?

hartnn (hartnn):

you will need to find 2 numbers with sum = +3 and product = -4

OpenStudy (anonymous):

1 + 2. 3 + -7=-4

hartnn (hartnn):

same 2 numbers should give sum = 3 and product of same 2 numbers should be -4 try to find such 2 numbers

OpenStudy (anonymous):

I can't figure it out

OpenStudy (cwrw238):

the 2 numbers are 4 and -1 because 4 * -1 = -4 and -1 + 4 = +3

OpenStudy (anonymous):

Thank you. It's 4 x -1= -4 -1 + 4= 3

OpenStudy (anonymous):

Is this the answer

OpenStudy (cwrw238):

no these 2 numbers go into 2 brackets as follows (t + 4)(t - 1) = 0

OpenStudy (cwrw238):

now either t +4 = 0 or t - 1 = 0 so there are 2 values of t -4 and 1

OpenStudy (cwrw238):

but going back to the original question t = x^2 so x^2 = 1 or -4